# Decomposable Submodular Maximization in Federated Setting

### v1

@@ 3 Federated Continuous Greedy
In this section we propose our Federated Continuous Greedy (FedCG) method. We start with a simplistic scenario of federated model with full participation which already shows some of the challenges that we have to overcome before delving into the partial participation model which is more computationally feasible.
@@
Bit complexity and accuracy trade-off.
For every client $i$ and every $\mathbf{x}\in\mathcal{P}$, the vector $\mathbf{v}_{i}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$ is determined by maximizing a linear function $\langle\mathbf{w},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$ over $\mathcal{P}$. In the case of a matroid polytope $\mathcal{P}$, this problem can be solved very efficiently. We can assume that $\mathbf{v}_{i}$ is a vertex of $\mathcal{P}$ and furthermore, since $\nabla\widehat{f_{i}}$ is a nonnegative vector, that this vertex corresponds to a base of matroid $\mathcal{M}$. Hence, without loss of generality $\mathbf{v}_{i}$ is the indicator vector of a base with $r$ ones and $n-r$ zeros. Thus it can be encoded using $O(r\log(n))$ bits, sublinear in the size of the ground set. On the other hand, the vector $\nabla\widehat{f_{i}}(\mathbf{x})$ itself requires $\tilde{O}(n)$ bits for encoding. In what follows we will see how restricting the bit complexity effects the accuracy, and the amount of computation that server should do.
FedCG with full participation.
First consider the case where clients can send their gradients. We proceeds in rounds. Initially $\mathbf{x}^{(0)}=0$. On the $t$-th round, first, the central server broadcasts the latest model $\mathbf{x}^{(t)}$ to all clients. Each client $i$ after receiving the update sets $\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$ and computes $\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})$. The server then aggregates local information via SecAgg, and computes $\nabla\widehat{F}(\mathbf{x}^{(t)})=\sum_{i=1}^{N}p_{i}\nabla\widehat{f}_{i}({\mathbf{x}^{(t)}})$. After receiving $\nabla\widehat{F}(\mathbf{x}^{(t)})$, the server computes $\mathbf{v}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ by maximizing a linear function subject to the matroid constraint and produces the new global model with learning rate $\eta$:
$\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\mathbf{v}^{(t)}$ (3)
It is clear that, similar to the centeralized CG, for large enough $T$ we get $F(\mathbf{x}^{(T)})\geq(1-\nicefrac{{1}}{{\mathrm{e}}})\texttt{OPT}$. Note that this simple framework has an advantage over centralized CG, it is taking advantage of the computational resources available at each client.
Second consider the more challenging case where clients can send at most $\tilde{O}(r)$ bits information. We see how this restriction effects the accuracy. Our algorithm, FedCG, proceeds in rounds. Initially $\mathbf{x}^{(0)}=0$. On the $t$-th round, first, the central server broadcasts the latest model $\mathbf{x}^{(t)}$ to all clients. Each client $i$ after receiving the update sets $\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$ and performs one step of continuous greedy approach to find a direction that best aligns with her local gradient:
@@
Lastly, clients send their update directions $\mathbf{v}_{1}^{(t)},\dots,\mathbf{v}_{N}^{(t)}$ to the secure aggregator to compute $\Delta^{(t)}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)}$. After receiving $\Delta^{(t)}$, the server produces the new global model with learning rate $\eta$:
@@
Even in this unrealistic setting where all clients participate in each round the convergence analysis requires new insights. In order to provide an approximation guarantee for our algorithm we shall obtain a lower bound on the function value improvement by taking the direction $\Delta^{(t)}$, that is providing a lower bound for $\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. However, each $\mathbf{v}_{i}^{(t)}$ is the projection of the local gradient into the matroid polytope and it does not carry information about the magnitude of the expected marginal contributions i.e. $\|\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\|$. Without assuming an assumption on the heterogeneity of local functions one can construct examples where a single element and corresponding client’s marginal contribution are significantly more dominant than others and hence taking direction $\Delta^{(t)}$ results in a very bad approximation guarantee.
@@
Note that monotonicity of $f_{i}$ implies that for every $\mathbf{x}\leq\mathbf{y}$ corrdinate-wise, it holds that $\widehat{f}_{i}(\mathbf{x})\leq\widehat{f}_{i}(\mathbf{x})$. Additionally, gradients are antitone i.e., for every $\mathbf{x}\leq\mathbf{y}$ corrdinate-wise, it holds that $\nabla\widehat{f}_{i}(\mathbf{x})\geq\nabla\widehat{f}_{i}(\mathbf{y})$. Thus as the algorithm advances and $t$ grows, $\gamma_{t}$’s change but never exceed the upper bound in (6). Additionally, in numerous instances, $\max_{t\in[T]}\gamma_{t}$ is relatively small. For instance, in Max Coverage problem each $f_{i}(S)$ is either 0 or 1 , depending if a client is covered by $S$ or not, thus for this problem $\gamma=1$.
@@ 3.1 Partial participation and client selection
$\bullet$ Small models and redundancy. It is often the case that FL models are small because of clients limited computational power or memory, it therefore is unnecessary to train an FL model on billions of clients. Note that for optimization problem (1) in many practical scenarios we are dealing with small models in comparison to the number of clients, this is because of the matroid constraint or simply because the size of the ground set is much smaller than the number of clients.
Here we discusse our sampling scheme which crucially does not violate clients’ privacy. The overall algorithm is given in Algorithm 1.
Unbiased Client Sampling Scheme.
At each communication round the server chooses an active client from $i\in[N]$ with probability $p_{i}$, and repeats this process $K$ times to obtain a multiset $A^{(t)}$ of size $K$ which may contain a client more than once. Then the aggregation step is $\Delta^{(t)}=\frac{1}{K}\sum\limits_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)}$ where $\mathbf{v}_{i}^{(t)}$ defined in (4).
@@
The following lemma allows us to bound the variance which in turn helps to provide our convergence guarantees. To show the following result, it is required to upper bound the difference between the improvement on the function value by taking the direction suggested by a selected client versus taking the direction obtained by averaging all the directions from the clients. That is bounding $|\langle\mathbf{v}_{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|$, for selected client $s$ in $A^{(t)}$. This in turn needs providing an upper bound for $\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{v}_{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$, for all $i\neq s$. At the heart, our proof relies on the properties of multilinear extensions of local submodular functions, the fact that each $\mathbf{v}_{i}^{(t)}$ corresponds to a base of the matroid, and Assumption 3.1.
@@ 4 Practical Federated Continuous Greedy
In convergence analyses of our algorithm there are two sources of randomness, one in the sampling schemes for client selection and the other in data sampling at each clients local data to estimate the gradients. In Algorithm 2 there are $\frac{T}{\tau}$ communications rounds between the server and clients. We define $\mathcal{I}_{\tau}=\{\tau i\mid i=1,2,3,\dots\}$ to denote the set of communication rounds with the server. Similar to Lemma 3.3:
Lemma 4.1 (Unbiased sampling scheme).
$\left\|\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j+1)})-\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j)})\right\|\leq\frac{m_{f_{i}}\sqrt{r}}{\tau}\left\|\widetilde{\mathbf{v}}_{i}^{(t,j)}\right\|\leq\frac{m_{f_{i}}r}{\tau}$
Observe that $L_{t}$ is upper bounded by $\max_{i}m_{f_{i}}\sqrt{r}$.
Lemma 4.2 (Bounded variance).
@@
[…]
@@
$\left(1-1/\mathrm{e}\right)\texttt{OPT}\leq\widehat{F}(\mathbf{x}^{(T)})+\frac{\tau r^{2}m_{F}}{2T}+$
$\frac{rD\tau}{T}+2\sigma r(1+\frac{\sqrt{\tau}}{\sqrt{KT\delta}})+\frac{2r^{1.5}Q\tau}{T}+\frac{\sqrt{\tau}(6Dr+2r^{1.5}Q)}{\sqrt{TK\delta}}$
5 Discrete problems in federated setting
Algorithm 3 Discrete Federated Submodular Maximization
@@ 5 Discrete problems in federated setting
13:     SecAgg: $\Delta^{(t)}$ (the sum of $\Delta_{i}$ it has received in this round.)
14:     Server updates: $S\leftarrow S\cup\{\operatornamewithlimits{argmax}_{e\in E\setminus S}\Delta^{(t)}[e]\}$
@@
While parallel SGD and continuous methods such as ours in this paper are commonly used as the main tool in federated optimization, we introduce a rather discrete approach to the field and believe it will find further applications. Our approach is inspired by recent works of Rafiey and Yoshida (2022); Kenneth and Krauthgamer (2023) on a seemingly unrelated topic. Rafiey and Yoshida (2022) introduced a method to sparsify a sum of submodular functions in a centralized setting which was improved by (Kenneth and Krauthgamer, 2023). We tailor their approach to the federated setting and discuss its effectiveness for discrete problems such as Facility Location and Maximum Coverage problems. Recall the optimization problem:
$\max_{S\in\mathcal{I}}\left\{F(S)=\sum_{i=1}^{N}f_{i}(S)\right\}$
where $\mathcal{M}=(E,\mathcal{I})$ is a matroid of rank $r$. At the heart of this approach is for clients to know their “importance” without sharing sensitive information. In the monotone case, the importance factor for client $i$ is defined as
@@
Let $\kappa=\tilde{O}(rn/\varepsilon^{2})$. Then each client sends its update in a randomized response manner, that is, the $i$-th client with probability $\kappa_{i}=\min\{1,\kappa\cdot w_{i}\}$ sends the scaled vector $(\frac{1}{\kappa_{i}})\Delta_{i}$ to the secure aggregator, and with the complement probability does not send anything. The secure aggregator computes $\Delta^{(t)}$; the sum of the update vectors it has received. The server updates set $S$:
$S\leftarrow S\cup\{\operatornamewithlimits{argmax}_{e:S\cup\{e\}\in\mathcal{M}}\Delta^{(t)}[e]\}$ (10)
The following theorem follows from the results in (Rafiey and Yoshida, 2022; Kenneth and Krauthgamer, 2023).
@@
For every $\varepsilon\in(0,1)$, Algorithm 3, with probability at least $1-1/n$, returns a subset $S\in\mathcal{I}$ such that $(1-\frac{1}{\mathrm{e}}-\varepsilon)\texttt{OPT}\leq F(S).$ Moreover, at each round $\tilde{O}(kn^{2}/\varepsilon^{2})$ clients participate.
5.1 Discrete examples
5.1.1 Facility Location Problem
For each client $i$ the importance factor is $w_{i}=\max\limits_{j\in E}\frac{c(i,j)}{F(\{j\})}$.
It is straightforward to see each client can compute its corresponding importance factor in a federated setting and using a secure aggregator without sharing its data with other clients. Each client $i$ sends a vector $(c(i,1),\dots,c(i,n))$ to the server and by simply summing up these vectors the server has a histogram over facilities. This histogram is then broadcasts to the clients where they can compute their own importance factor (see Algorithm 4). This requires only two communication rounds. Having $w_{i}$ on hand we can execute Algorithm 3 for Max Facility Location problem. Note that at each round of Algorithm 3 the expected number of clients participating is $\tilde{O}(kn^{2}/\varepsilon^{2})$, independent of $N$.
5.1.2 Maximum Coverage Problem
Similar to the case of Facility Location the importance factors can be computed in federated setting by having the membership histograms over the groups (see Algorithm 5). Applying Algorithm 3 yields $(1-1/e-\varepsilon)$ approximation guarantee where at each round the expected number of participating clients is $\tilde{O}(kn^{2}/\varepsilon^{2})$, independent of $N$.

### v2

@@ 3 Federated Continuous Greedy
In this section, we propose our Federated Continuous Greedy (FedCG) method. We start with a simplistic scenario of federated model with full participation which already shows some of the challenges that we have to overcome before delving into the partial participation model which is more computationally feasible.
@@
Bit complexity and accuracy trade-off. For every client $i$ and every $\mathbf{x}\in\mathcal{P}$, the vector $\mathbf{v}_{i}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$ is determined by maximizing a linear function $\langle\mathbf{w},\nabla\widehat{f_{i}}(\mathbf{x})\rangle$ over matroid polytope $\mathcal{P}$. This problem can be solved very efficiently. We can assume that $\mathbf{v}_{i}$ is a vertex of $\mathcal{P}$ and furthermore, since $\nabla\widehat{f_{i}}$ is a nonnegative vector, that this vertex corresponds to a base of matroid $\mathcal{M}$. Hence, without loss of generality $\mathbf{v}_{i}$ is the indicator vector of a base with $r$ ones and $n-r$ zeros. Thus it can be encoded using $O(r\log(n))$ bits, sublinear in the size of the ground set. On the other hand, the vector $\nabla\widehat{f_{i}}(\mathbf{x})$ itself requires $\tilde{O}(n)$ bits for encoding. In what follows we will see how restricting the bit complexity effects the accuracy, and the amount of computation that server should do.
FedCG with full participation. First consider the case where clients can send their gradients. We proceeds in rounds. Initially $\mathbf{x}^{(0)}=0$. On the $t$-th round, first, the central server broadcasts the latest model $\mathbf{x}^{(t)}$ to all clients. Each client $i$ after receiving the update sets $\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$ and computes $\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})$. The server then aggregates local information via SecAgg, and computes $\nabla\widehat{F}(\mathbf{x}^{(t)})=\sum_{i=1}^{N}p_{i}\nabla\widehat{f}_{i}({\mathbf{x}^{(t)}})$. After receiving $\nabla\widehat{F}(\mathbf{x}^{(t)})$, the server computes $\mathbf{v}^{(t)}=\operatornamewithlimits{argmax}_{\mathbf{w}\in\mathcal{P}}\langle\mathbf{w},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$ by maximizing a linear function subject to the matroid constraint and produces the new global model with learning rate $\eta$: $\mathbf{x}^{(t+1)}\leftarrow\mathbf{x}^{(t)}+\eta\mathbf{v}^{(t)}$. It is clear that, similar to the centeralized CG, large enough $T$ yields $F(\mathbf{x}^{(T)})\geq(1-\nicefrac{{1}}{{\mathrm{e}}})\texttt{OPT}$. This simple framework has an advantage over centralized CG, it is taking advantage of the computational resources available at each client.
Second consider a more challenging case where clients can send at most $\tilde{O}(r)$ bits information. We see how this restriction effects the accuracy. Our algorithm, FedCG, proceeds in rounds. Initially $\mathbf{x}^{(0)}=0$. On the $t$-th round, first, the central server broadcasts the latest model $\mathbf{x}^{(t)}$ to all clients. Each client $i$ after receiving the update sets $\mathbf{x}_{i}^{(t)}=\mathbf{x}^{(t)}$ and performs one step of continuous greedy approach to find a direction that best aligns with her local gradient:
@@
Lastly, clients send their update directions $\mathbf{v}_{1}^{(t)},\dots,\mathbf{v}_{N}^{(t)}$ to the secure aggregator to compute $\Delta^{(t)}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{i}^{(t)}$. After computing $\Delta^{(t)}$, the server produces the new global model with learning rate $\eta$:
@@
Even in this unrealistic setting where all clients participate in each round the convergence analysis requires new insights. In order to provide an approximation guarantee for our algorithm, we shall obtain a lower bound on the function value improvement by taking the direction $\Delta^{(t)}$, that is providing a lower bound for $\langle\Delta^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle$. However, each $\mathbf{v}_{i}^{(t)}$ is the projection of the local gradient into the matroid polytope and it does not carry information about the magnitude of the expected marginal contributions i.e. $\|\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\|$. Without assuming an assumption on the heterogeneity of local functions one can construct examples where a single element and corresponding client’s marginal contribution are significantly more dominant than others and hence taking direction $\Delta^{(t)}$ results in a very bad approximation guarantee.
@@
Monotonicity of $f_{i}$ implies that for every $\mathbf{x}\leq\mathbf{y}$ coordinate-wise, it holds that $\widehat{f}_{i}(\mathbf{x})\leq\widehat{f}_{i}(\mathbf{x})$. Additionally, gradients are antitone i.e., for every $\mathbf{x}\leq\mathbf{y}$ coordinate-wise, it holds that $\nabla\widehat{f}_{i}(\mathbf{x})\geq\nabla\widehat{f}_{i}(\mathbf{y})$. Thus as the algorithm advances and $t$ grows, $\gamma_{t}$’s change but never exceed the upper bound in (5). Additionally, in numerous instances, $\max_{t\in[T]}\gamma_{t}$ is relatively small. For instance, in Max Coverage problem each $f_{i}(S)$ is either 0 or 1 , depending if a client is covered by $S$ or not, thus for this problem $\gamma=1$.
@@ 3.1 Partial Participation and Client Selection
$\bullet$ Small models and redundancy. It is often the case that FL models are small because of clients limited computational power or memory, it therefore is unnecessary to train an FL model on billions of clients. Note that for optimization problem (1) in many practical scenarios models have smaller size in comparison to the number of clients, this is because of the matroid constraint or simply because the size of the ground set is much smaller than the number of clients.
Here we discuss our sampling scheme which crucially does not violate clients’ privacy; see Algorithm 1.
Unbiased client sampling scheme. At each communication round the server chooses an active client from $i\in[N]$ with probability $p_{i}$, and repeats this process $K$ times to obtain a multiset $A^{(t)}$ of size $K$ which may contain a client more than once. Then the aggregation step is $\Delta^{(t)}=\frac{1}{K}\sum_{i\in A^{(t)}}\mathbf{v}_{i}^{(t)}$ where $\mathbf{v}_{i}^{(t)}$ defined in (3).
@@
The following lemma allows us to bound the variance which in turn helps to provide our convergence guarantees. To show the following result, it is required to upper bound the difference between the improvement on the function value by taking the direction suggested by a selected client versus taking the direction obtained by averaging all the directions from the clients. That is bounding $|\langle\mathbf{v}_{s},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle-\langle\sum_{i=1}^{N}\mathbf{v}_{i}^{(t)},\nabla\widehat{F}(\mathbf{x}^{(t)})\rangle|$, for a selected client $s$ in $A^{(t)}$. This in turn needs providing an upper bound for $\langle\mathbf{v}_{i}^{(t)},\nabla\widehat{f_{i}}(\mathbf{x}^{(t)})\rangle-\langle\mathbf{v}_{s},\nabla\widehat{f}_{i}(\mathbf{x}^{(t)})\rangle$, for all $i\neq s$. At the heart, our proof relies on the properties of multilinear extensions of local submodular functions, the fact that each $\mathbf{v}_{i}^{(t)}$ corresponds to a base of the matroid, and Assumption 3.1.
@@ 4 Practical Federated Continuous Greedy
In convergence analyses of our algorithm there are two sources of randomness, one in the sampling schemes for client selection and the other in data sampling at each clients local data to estimate the gradients. In Algorithm 2 there are $\frac{T}{\tau}$ communications rounds between the server and clients. Define $\mathcal{I}_{\tau}=\{\tau i\mid i=1,2,3,\dots\}$ to denote the set of communication rounds with the server. Similar to Lemma 3.3:
Lemma 4.1 (Unbiased sampling scheme).
$\left\|\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j+1)})-\nabla\widehat{f}_{i}(\mathbf{x}_{i}^{(t,j)})\right\|$ $\leq\frac{m_{f_{i}}\sqrt{r}}{\tau}\left\|\widetilde{\mathbf{v}}_{i}^{(t,j)}\right\|$
$\leq\frac{m_{f_{i}}r}{\tau}$
Observe that $L_{t}$ is upper bounded by $\max_{i}m_{f_{i}}\sqrt{r}$.
Lemma 4.2 (Bounded variance).
@@
$(1-(1-\eta)^{T/\tau})\texttt{OPT}$
$\quad\quad\leq\widehat{F}(\mathbf{x}^{(T/\tau)})+\frac{T\eta^{2}r^{2}m_{F}}{2\tau}$
$\quad\quad+(\overbrace{\eta r\sum_{t=1}^{T/\tau}\gamma_{t}}^{\text{heterogeneity}}+\overbrace{2\sigma r+2\eta r^{1.5}\sum_{t=1}^{T/\tau}L_{t})}^{\text{local steps}}$
$\quad\quad+\overbrace{{\sqrt{T}(6\eta r\sum_{t=1}^{T/\tau}\gamma_{t}+2\sigma r+2\eta r^{1.5}\sum_{t=1}^{T/\tau}L_{t}))}/{\sqrt{K\tau\delta}}}^{\text{client sampling}})$
@@
$\left(1-1/\mathrm{e}\right)\texttt{OPT}$ $\leq\widehat{F}(\mathbf{x}^{(T)})+\frac{\tau r^{2}m_{F}}{2T}$
$+\frac{rD\tau}{T}+2\sigma r(1+\frac{\sqrt{\tau}}{\sqrt{KT\delta}})+\frac{2r^{1.5}Q\tau}{T}$
$+\frac{\sqrt{\tau}(6Dr+2r^{1.5}Q)}{\sqrt{TK\delta}}$
Remark 4.4.
From a fractional solution $\mathbf{x}^{(T)}\in\mathcal{P}$ returned by Algorithms 1 and 2 one can obtain a solution for (1) using appropriate rounding schemes. Rounding schemes such as pipage rounding (Călinescu et al., 2011), swap rounding (Chekuri et al., 2010), or greedy rounding are oblivious; they do not require access to the objective function. Therefore, the server can utilize these rounding schemes without needing to know the decomposable function itself.
5 Discrete Algorithm in Federated Setting
Algorithm 3 Federated Discrete Greedy
@@ 5 Discrete Algorithm in Federated Setting
13:     SecAgg: $\Delta^{(t)}$ (the sum of $\Delta_{i}$ received in this round.)
14:     Server updates: $S\leftarrow S\cup\{\operatornamewithlimits{argmax}\limits_{e:S\cup\{e\}\in\mathcal{M}}\Delta^{(t)}[e]\}$
@@
While parallel SGD and continuous methods such as ours in this paper are commonly used as the main tool in federated optimization, we introduce a rather discrete approach to the field and believe it will find further applications. Our approach is inspired by recent works of Rafiey & Yoshida (2022); Kenneth & Krauthgamer (2023) on a seemingly unrelated topic. Rafiey & Yoshida (2022) introduced a method to sparsify a sum of submodular functions in a centralized setting which was improved by (Kenneth & Krauthgamer, 2023). We tailor their approach to the federated setting and discuss its effectiveness for discrete problems such as Facility Location and Maximum Coverage problems. At the heart of this approach is for clients to know their “importance” without sharing sensitive information. In the monotone case, the importance factor for client $i$ is defined as:
@@
Let $\kappa=\tilde{O}(rn/\varepsilon^{2})$. Then each client sends its update in a randomized response manner, that is, the $i$-th client with probability $\kappa_{i}=\min\{1,\kappa\cdot w_{i}\}$ sends the scaled vector $(\frac{1}{\kappa_{i}})\Delta_{i}$ to the secure aggregator, and with the complement probability does not send anything. The secure aggregator computes $\Delta^{(t)}$; the sum of the update vectors it has received. The server updates $S\leftarrow S\cup\{\operatornamewithlimits{argmax}_{e:S\cup\{e\}\in\mathcal{M}}\Delta^{(t)}[e]\}$. The following theorem follows from the sparsification results in (Rafiey & Yoshida, 2022; Kenneth & Krauthgamer, 2023) and approximation guarantees of the greedy algorithm (Nemhauser et al., 1978).
@@
For every $\varepsilon\in(0,1)$, Algorithm 3, with probability at least $1-1/n$, returns a subset $S\in\mathcal{I}$ such that $({1}/{2}-\varepsilon)\texttt{OPT}\leq F(S).$
Under a cardinality constraint (uniform matroid), Algorithm 3, with probability at least $1-1/n$, returns a subset $S\in\mathcal{I}$ such that $(1-{1}/{\mathrm{e}}-\varepsilon)\texttt{OPT}\leq F(S).$
Moreover, in expectation, at each round $\tilde{O}(rn^{2}/\varepsilon^{2})$ clients participate where $r$ is the rank of the matroid.
Facility Location Problem. Let $\mathcal{C}$ be a set of $N$ clients and $E$ be a set of facilities with $|E|=n$. For $c:\mathcal{C}\times E\to\mathbb{R}$ let the $i$-th client’s score function over a subset of facilities be $f_{i}(A)=\max_{j\in A}c(i,j)$. The objective for Max Facility Location is $\max_{S\subseteq E,|S|\leq k}\sum_{i=1}^{N}\max_{j\in S}c(i,j)$. For each client $i$ the importance factor is $w_{i}=\max_{j\in E}\frac{c(i,j)}{F(\{j\})}$.
In several situations computing the importance factors is straightforward. For instance, in movie recommendation systems where a facility location objective is used (see Appendix D), the average rating i.e., $F(\{j\})$, for all movies are publicly available. Hence, each client knows its own importance factor. Having $w_{i}$, we can apply Algorithm 3. Doing so, note that at each round the expected number of clients participating is $\tilde{O}(kn^{2}/\varepsilon^{2})$, independent of $N$.
In other situations, clients can compute their importance factors in a federated setting, using a secure aggregator and without sharing their data with other clients. Further details on this and discussions about the Max Coverage problem are presented in Appendix D.